# SReferences

### v1

@@ Abstract
Fusion-based quantum computation is a promising quantum computing model where small-sized photonic resource states are simultaneously entangled and measured by fusion gates. Such operations can be readily implemented with scalable photonic hardware: resource states can be deterministically generated by quantum emitters and fusions require only shallow linear-optical circuits. Here, we propose fusion-based architectures tailored to the capabilities and noise models in quantum emitters. We show that high tolerance to dominant physical error mechanisms can be achieved, with fault-tolerance thresholds of 8% for photon loss, 4% for photon distinguishability between emitters, and spin noise thresholds well below coherence times for typical spin-photon interfaces. Our construction and analysis provide guidelines for the development of photonic quantum hardware targeting fault-tolerant applications with quantum emitters.
@@ I Introduction
Figure 1: Decomposition of FFCC lattice. (a) Cluster state for FFCC lattice. In the measurement-based picture, all qubits are measured in the Pauli-$X$ basis to provide syndrome data for error correction. (b) The lattice can be decomposed into a fusion network consisting of six linear cluster states localized in one unit cell, with fusions applied along the horizontal edges. (c) Original FFCC fusion network from Ref. [11] where resource states (identified by different colors) extend into other unit cells and fusions connect different layers. (d) Linear-optic circuit for a type-II $\{XX,ZZ\}$ fusion and possible outcomes based on different detection patterns. Successful fusions project two input qubits into a Bell state |$\ket{\psi^{\pm}}$, while fusion failure leads to single-qubit measurements. (e) Edge split rule is used to decompose the FFCC lattice into (c). (f) Node split rule is applied to convert a qubit into fusions between two additional qubits, resulting in (b).
Here, we present a fusion-based architecture tailored to quantum emitters and analyze how relevant physical processes affect its fault-tolerant performance. The architecture is based on the foliated Floquet color code (FFCC) [11] where each resource state can be deterministically generated by a single quantum emitter, with high tolerance to photon loss and qubit errors. We develop a framework to describe how the dominant noise processes in quantum emitters propagate to the logical qubit level of the architecture and analyze its performance in terms of fault-tolerant thresholds to physical noises. Specifically, we investigate photon distinguishability, Markovian spin noise in spin-photon interfaces, and photon loss. We show that our tailored architectures are highly resilient to these noises, as summarized in Table. 1.
@@ II FFCC lattice and fusion network
In Ref. [11], the authors introduced the foliated Floquet color code (FFCC) lattice, constructed by foliating the Floquet color code [19, 20]. The code can be understood both in the circuit based model and by looking at the graph state version of it in Fig. 1a. Here we describe it in terms of the latter. In general, a graph state is associated with a certain graph, where each vertex represents a qubit prepared in the state $\ket{+}$ and edges represent $CZ$ gates. In the Floquet color code, data qubits are situated at the vertices of a hexagonal lattice, and check operations are determined by a coloration of the lattice. All faces are assigned a color, such that no two adjacent faces share the same color. Edge colors are determined by the color of the faces at either end of an edge, e.g. a green edge connects two green faces. The check operators of the Floquet color code are built up from two-body check measurements defined by the colored edges [20].
@@ II.1 Fusion-based quantum computation
Here we are interested in optical implementations of fusion networks. Photonic qubits can be represented by two orthogonal modes encoded, e.g., in polarization, frequency, time-bin and photon number. Dual-rail encoding, which uses the presence of a photon in either of two orthogonal modes as a qubit, offers loss-tolerance properties which make it appealing for FBQC [2]. With linear optics, fusion measurements can be implemented on dual-rail qubits by a simple four-mode circuit, with photon counting detectors on each mode (Fig. 1d). This linear circuit performs fusion measurement with 50% success probability. When the fusion fails, measuring single qubit operators $Z_{a}$ and $Z_{b}$ retrieves the $Z_{a}Z_{b}$ outcome. Single qubit gates can be used to rotate the bases of these measurements, for example, a Hadamard on one of the input qubits results in an $\{XZ,ZX\}$ fusion, and Hadamards on both inputs results in a measurement that performs a $\{ZZ,XX\}$ measurement, but retrieves $XX$ on fusion failure.
Fusion networks can be conveniently described in the stabilizer formalism [22], where quantum states are represented as eigenstates with eigenvalue $+1$ of a set of operators, e.g., the graph state in Fig. 1a is the eigenstate of a set of operators $X_{i}\prod_{j\in N(i)}Z_{j}$, where $i$ is the qubit index, and $N(i)$ is the neigborhood of qubit $i$ which refer to all qubits connected to it by an edge in the graph. In the fusion-based picture, fusion measurements and resource states can both be represented by subgroups of the Pauli group, denoted $\mathcal{F}$ and $\mathcal{R}$, respectively. $\mathcal{F}$ is the set of operators measured by all fusions, and $\mathcal{R}$ is the stabilizer group of all resource states [2]. The intersection of these two groups gives the surviving stabilizer group $C:=\mathcal{R}\cap\mathcal{F}$ [2], a group of check operators that return a $+1$ eigenvalue in the absence of errors. Errors in the resource states or fusion measurements that anti-commute with elements of $C$ result in negative eigenvalues for one or more check operators. The set of eigenvalues of check operators is called the syndrome, and is interpreted by a decoder for error correction in FTQC.
@@ II.2 Constructing fusion networks from lattice decomposition
An edge split rule is constructed by noting that an $\left\{XZ,ZX\right\}$ fusion is equivalent to a $CZ$ gate between the qubits involved in the fusion, before both qubits are measured in the $X$ basis. For any edge in a cluster state, we can apply an edge split, introducing a fusion measurement, and assert that an $\{XZ,ZX\}$ fusion will recover the same stabilizer group as the cluster state protocol, up to Pauli frame updates due to the measurement outcomes. A physical qubit can only participate in a single fusion, but introducing a redundantly-encoded qubit at a lattice site allows for many adjacent edge-splits, as shown schematically in Fig. 1e.
@@
For node splits, an equivalence is identified between the measurement of a cluster state node in the $X$ basis and an $\left\{XX,ZZ\right\}$ fusion between two nodes that are introduced at the same site. One of the new nodes is connected to a subset of the neighbours of the original node, with the other connected to the complement, as shown in Fig. 1f. A similar concept of node splitting was introduced in Ref. [23] but not in the context of fusions.
@@
These operations can be systematically applied to the cluster state until target resource states are obtained. Any remaining qubits that do not take part in fusion gates have implicit $X$ measurements on them, and so we follow Refs. [2, 11] and directly prepare the resource states where these qubits have been projected onto $\ket{+}$.
@@
Applying a node split on each check qubit of the FFCC lattice (white circles in Fig. 1a) replaces it with two qubits, which are measured by an $\{XX,ZZ\}$ fusion. The two qubits (black) that were connected to the check qubit prior to the node split do not participate in fusions. Therefore, by measuring these qubits in $X$-basis and projecting them into $\ket{+}$, we obtain the target resource state—linear cluster state  [^1] Up to Hadamard gates on the remaining qubits, which when compiled into fusions do not affect the measurement bases.. This results in a fusion network consisting of linear cluster states shown in Fig. 1b, which we refer to as “synchronous FFCC”, where qubits on the same layer can be simultaneously generated. As we shall describe later, this synchronicity allows for the use of RUS fusions, providing significant practical advantages for the implementation and noise tolerance of this scheme. As a comparison, the original FFCC fusion architecture (Fig. 1c) was constructed using the edge split rule.
Figure 2: Fusion network for synchronous FFCC. Resource states (RS) are generated by an array of quantum emitters (red circles). Fusions (yellow) are performed between photons to obtain pair-wise parity checks (cyan wavy lines) in layers.
@@
Another important advantage of the synchronous FFCC network is that resource states remain spatially local and interact with only three neighbouring states. For resource states generated by quantum emitters, as described in the next section, this feature ensures that routing requirements for photons from different emitters remain constant with the size of the computation (Fig. 2), in contrast to the original scheme.
@@ III Resource state generator
An encoded linear chain state can also be generated by a single emitter, where each qubit is redundantly encoded in $m$ photons [17]. As shown in Fig. 3b, this is done by omitting some interstitial Hadamard gates on the emitter in Step 2b of the circuit, which generates a linear chain with each encoded qubit (grey dashed circle) corresponding to an $m$-photon GHZ state [17]. In the graph state picture these additional photons can be described as branches on the linear chain. This encoding is essential for improving the loss and fault tolerances of the fusion-based architecture, as we shall cover in Sec. VI.
@@ V.2 Spin errors during resource state generation
In the simulation, we sample spin errors from the Pauli group $\mathcal{P}=\{I,X,Z,Y\}$ at each step during state generation with probabilities $\{1-p^{(s)},\frac{p^{(s)}}{3},\frac{p^{(s)}}{3},\frac{p^{(s)}}{3}\}$, where $p^{(s)}$ is the depolarizing error probability. Therefore, we have $\frac{p^{(s)}}{3}=\frac{1}{4}\big{[}1-\exp(-\tau/T_{2})\big{]}$ [18] where $\tau$ is the time required to generate a photon. With a depolarizing channel, $T_{2}=T_{1}$, allowing us to infer a lower bound on the threshold for $T_{2}$ from $p^{(s)}$ without further assumptions.
@@ VI.1 Logical Fusion with repetition codes
For synchronous FFCC, we seek to perform logical $\left\{XX,ZZ\right\}$ fusions on qubits encoded in a $Z$-repetition code of size $m$. In order to recover the $\overline{XX}$ outcome, all $m$ physical $XX$ outcome must be obtained (Eq. (2)), but for $\overline{ZZ}$ a single physical $ZZ$ outcome is sufficient. As such, it is beneficial to perform rotated physical fusions on each pair of physical qubits, which measure $XX$ and $ZZ$ on success, and recover the $XX$ outcome upon fusion failure (Sec. II.1). In this manner, the $\overline{XX}$ outcome is protected by the choice of physical fusion, while the $\overline{ZZ}$ outcome is protected by redundancy of the repetition code.
The probabilities of errors and erasures on fusion outcomes using a repetition code with fixed code size (REP) are calculated analytically in Supplementary Note I. We employ a simple majority-voting decoding strategy for errors on the $k$ non-erased $ZZ$ outcomes, such that $\left\lfloor\frac{k-1}{2}\right\rfloor$ errors can be corrected. The $Z$-repetition code offers no tolerance to $Z$ error, so the error rate on the $\overline{XX}$ outcome, provided it is not erased, is the probability of having an odd number of erroneous $XX$ outcomes.
@@ VI.2 Repeat-until-success logical fusions
The probabilities for different logical fusion outcomes of a single RUS fusion under photon loss and errors can be analytically derived, see Supplementary Note I. Using these analytical expressions, logical fusion outcomes are sampled to obtain a syndrome graph for calculating fault-tolerant thresholds.
@@ VII Fault-tolerant Thresholds
The loss-tolerance thresholds are presented in Fig. 7a, for various code sizes $m$ for REP and maximal number of physical fusion attempts $N$ for RUS. The highest REP loss threshold obtained is $\ell=2.29\%$ for $m=5$. For RUS logical fusion, we observe that $\ell$ improves significantly as the number of maximally allowed fusion attempts $N$ increases, achieving $\ell=7.4\%$ with $N=10$. Using a strategy where additional $N$ re-attempts of RUS are allowed with both emitters re-initialized from the previous layer of logical fusions (Supplementary Note II), $\ell$ can reach up to $8\%$.
@@
Fig. 8a plots the simulated spin $Z$ error threshold with the number of fusion attempts per logical fusion, with $p^{(s)}_{X}=p^{(s)}_{Y}=0$. We observe that for spin errors occurred during the generation of encoded linear chains, the RUS threshold increases up to $p^{(s)}_{Z}=0.58\%$ at $N=10$ whereas for REP it peaks at $p^{(s)}_{Z}=0.19\%$ then steadily decreases to $0.095\%$. As discussed in Sec. V.2, a spin $Z$ error will localize as a $Z$ error in the next emitted photon. For REP the number of fusion attempts $m$ is fixed, such that when $m$ grows the probability of flipping $\overline{XX}$ also increases, and this leads to the decreasing threshold for high $m$. In the limit of low $m$, however, more fusion attempts increases the fusion success probability resulting in a more connected and hence more robust lattice, with an optimal threshold at $m=5$. For RUS, most logical fusions terminate in a few attempts assuming no photon loss (Supplementary Note I), so that only $Z$ errors occurring in the photons emitted during those attempts contribute to a flip in $\overline{XX}$. As a consequence, the spin $Z$ error threshold saturates as we increase $N$.
For spin depolarizing error (Fig. 8b), the RUS (REP) threshold peaks at $p^{(s)}=0.2\%$ ($0.125\%$), and both gradually reduce with the number of fusion attempts. This can be understood by first considering the propagation of spin $X$/$Y$ errors during resource state generation into photon errors, before mapping these photon errors to logical fusion errors. A spin $X$/$Y$ error during the generation of first $k\leq m$ photons in a logical qubit will lead to $X/Y$ errors in the subsequently generated $m-k$ photons  [^2] More precisely, a spin $Y$ error results in a $Y$ error on the first emitted photon and $X$ errors on the next generated photons in the current logical qubit $i$, as well as a $Z$ error on the first photon of the next qubit $i+1$., as well as a $Z$ photon error on the first photon of the next logical qubit (Sec. V.2). This implies the probability of flipping the $ZZ$ outcome (hence $\overline{ZZ}$) increases with number of fusion attempts, since a larger $m$ results in a greater probability of $X/Y$ spin errors within an $m$-photon logical qubit. Both error thresholds therefore reduce at higher values of $m$/$N$. Moreover, for REP, majority voting becomes ineffective at correcting erroneous $ZZ$ outcome due to a high weight of $X/Y$ errors from a single spin error.

### v2

@@ Abstract
Fusion-based quantum computation is a promising quantum computing model where small-sized photonic resource states are simultaneously entangled and measured by fusion gates. Such operations can be readily implemented with scalable photonic hardware: resource states can be deterministically generated by quantum emitters and fusions require only shallow linear-optical circuits. Here, we propose fusion-based architectures tailored to the capabilities and noise models in quantum emitters. We show that high tolerance to dominant physical error mechanisms can be achieved, with fault-tolerance thresholds of 8% for photon loss, 4% for photon distinguishability between emitters, and spin noise thresholds well above memory-induced errors for typical spin-photon interfaces. Our construction and analysis provide guidelines for the development of photonic quantum hardware targeting fault-tolerant applications with quantum emitters.
@@ I Introduction
Figure 1: Decomposition of FFCC lattice. (a) Cluster state for FFCC lattice. In the measurement-based picture, all qubits are measured in the Pauli-$X$ basis to provide syndrome data for error correction. (b) The lattice is decomposed into a fusion network consisting of six linear cluster states localized in one unit cell, with fusions applied along the horizontal edges. (c) Original FFCC fusion network [11] where resource states (identified by different colors) extend into other unit cells and fusions connect different layers. (d) Linear-optic circuit for a type-II $\{XX,ZZ\}$ fusion and possible outcomes based on different detection patterns. Successful fusions project two input qubits into a Bell state |$\ket{\psi^{\pm}}$, while fusion failure leads to single-qubit measurements. (e) Edge split rule is used to decompose the FFCC lattice into (c). (f) Node split rule is applied to convert a qubit into fusions between two additional qubits. When combined with $X$ measurements on the neighbors, this results in (b).
Here, we present a fusion-based architecture tailored to quantum emitters and analyze how relevant physical processes affect its fault-tolerant performance. The architecture is based on the foliated Floquet color code (FFCC) [11] where each resource state can be deterministically generated by a single quantum emitter, with high tolerance to photon loss and qubit errors. We develop a framework to describe how the dominant noise processes in quantum emitters propagate to the logical qubit level of the architecture and analyze its performance in terms of fault-tolerant thresholds to physical noises. Specifically, we investigate photon distinguishability, Markovian spin noise in spin-photon interfaces, and photon loss. We show that our tailored architectures are highly resilient to these noises, as summarized in Table 1.
@@ II FFCC lattice and fusion network
In Ref. [11], the authors introduced the foliated Floquet color code (FFCC) lattice, constructed by foliating the Floquet color code [19, 20]. The code can be understood both in the circuit based model and by looking at the graph state version of it in Fig. 1a. Here we describe it in terms of the latter. In general, a graph state is associated with a certain graph, where each vertex represents a qubit prepared in the state $\ket{+}$ and edges represent $CZ$ gates. In the Floquet color code, data qubits are situated at the vertices of a hexagonal lattice, and check operations are determined by a coloration of the lattice. All faces are assigned a color, such that no two adjacent faces share the same color. Edge colors are determined by the color of the faces at either end of an edge, e.g., a green edge connects two green faces. The check operators of the Floquet color code are built up from two-body check measurements defined by the colored edges [20].
@@ II.1 Fusion-based quantum computation
Here we are interested in optical implementations of fusion networks. Photonic qubits can be represented by two orthogonal modes encoded, e.g., in polarization, frequency, time-bin and photon number. Dual-rail encoding, which uses the presence of a photon in either of two orthogonal modes as a qubit, offers loss-tolerance properties which make it appealing for FBQC [2]. With linear optics, fusion measurements can be implemented on dual-rail qubits by a simple four-mode circuit, with photon counting detectors on each mode (Fig. 1d). This linear circuit performs fusion measurement with 50% success probability. When the fusion fails, measuring single qubit operators $Z_{a}$ and $Z_{b}$ retrieves the $Z_{a}Z_{b}$ outcome. Single qubit gates can be used to rotate the bases of these measurements, e.g., a Hadamard on one of the input qubits results in a $\{XZ,ZX\}$ fusion, and Hadamards on both inputs lead to a $\{ZZ,XX\}$ fusion which retrieves $XX$ upon fusion failure.
Fusion networks can be conveniently described in the stabilizer formalism [22], where quantum states are represented as eigenstates with eigenvalue $+1$ of a set of operators, e.g., the graph state in Fig. 1a is the eigenstate of a set of operators $X_{i}\prod_{j\in\mathcal{N}_{i}}Z_{j}$, where $i$ is the qubit index, and $\mathcal{N}_{i}$ is the neigborhood of qubit $i$ which refer to all qubits connected to it by an edge in the graph. In the fusion-based picture, fusion measurements and resource states can both be represented by subgroups of the Pauli group, denoted $\mathcal{F}$ and $\mathcal{R}$, respectively. $\mathcal{F}$ is the set of operators measured by all fusions, and $\mathcal{R}$ is the stabilizer group of all resource states [2]. The intersection of these two groups gives the surviving stabilizer group $C:=\mathcal{R}\cap\mathcal{F}$ [2], a group of check operators that return a $+1$ eigenvalue in the absence of errors. Errors in the resource states or fusion measurements that anti-commute with elements of $C$ result in negative eigenvalues for one or more check operators. The set of eigenvalues of check operators is called the syndrome, and is interpreted by a decoder for error correction in FTQC.
@@ II.2 Constructing fusion networks from lattice decomposition
An edge split rule is constructed by noting that a $\left\{XZ,ZX\right\}$ fusion is equivalent to a $CZ$ gate between the qubits involved in the fusion, before both qubits are measured in the $X$ basis (Fig. 1e). For any edge in a cluster state, we can apply an edge split, introducing a fusion measurement, and assert that a $\{XZ,ZX\}$ fusion will recover the same stabilizer group as the cluster state protocol, up to Pauli frame updates due to the measurement outcomes.
@@
For node splits, an equivalence is identified between the measurement of a cluster state node in the $X$ basis and a $\left\{XX,ZZ\right\}$ fusion between two nodes that are introduced at the same site. One of the new nodes is connected to a subset of the neighbours of the original node, with the other connected to the complement, as shown in Fig. 1f. A similar concept of node splitting was introduced in Ref. [23] but not in the context of fusions.
@@
The proofs for both rules are provided in Supplementary Note I. In general, these decomposition rules can be systematically applied to any cluster state until target resource states are obtained. Any remaining qubits that do not take part in fusion gates have implicit $X$ measurements on them, and so we follow Refs. [2, 11] and directly prepare the resource states where these qubits have been projected onto $\ket{+}$.
@@
To obtain the synchronous FFCC fusion network in Fig. 1b from decomposition of the FFCC lattice, we first apply a node split on each check qubit of the lattice (white circles in Fig. 1a) and replaces it with two qubits, which are measured by a $\{XX,ZZ\}$ fusion. Next, since the two qubits (black) that were connected to the check qubit prior to the node split do not participate in fusions, we consider them as “virtual qubits” that are not generated by actual hardware, but are rather assumed to be measured in the $X$ basis resulting in $+1$ outcomes. $X$ measurements on these qubits remove them from the graph and have their neighbors connected, then add Hadamard rotations on the neighboring qubits which participate in the $\{XX,ZZ\}$ fusion (Supplementary Note I). As such, the fusion failure basis is switched to $XX$. This results in a fusion network consisting of linear cluster states shown in Fig. 1b, which we refer to as “synchronous FFCC”, where qubits on the same layer can be simultaneously generated. As we shall describe later, this synchronicity allows for the use of RUS fusions, providing significant practical advantages for the implementation and noise tolerance of this scheme. As a comparison, the original FFCC fusion architecture (Fig. 1c) was constructed using edge splits.
@@
Another important advantage of the synchronous FFCC network is that resource states remain spatially local and interact with only three neighbouring states. For resource states generated by quantum emitters, this feature ensures that routing requirements for photons from different emitters remain constant with the size of the computation (Fig. 2), in contrast to the original scheme.
Figure 2: Fusion network for synchronous FFCC. Resource states (RS) are generated by an array of quantum emitters (red circles). Fusions (yellow) are performed between photons to obtain pair-wise parity checks (cyan wavy lines) layers by layers.
@@ III Resource state generator
An encoded linear chain state can also be generated by a single emitter, where each qubit is redundantly encoded in $m$ photons [17]. As shown in Fig. 3b, this is done by omitting some interstitial Hadamard gates on the emitter in Step 2b of the circuit, which generates a linear chain with each encoded qubit (grey dashed circle) in an $m$-photon repetition code [17]. In the graph state picture these additional photons can be described as branches on the linear chain. This encoding is essential for improving the loss and fault tolerances of the fusion-based architecture, as we shall cover in Sec. VI.
@@ V.2 Spin errors during resource state generation
In the simulation, we sample spin errors from the Pauli group $\mathcal{P}=\{I,X,Z,Y\}$ at each step during state generation with probabilities $\{1-p^{(s)},\frac{p^{(s)}}{3},\frac{p^{(s)}}{3},\frac{p^{(s)}}{3}\}$, where $p^{(s)}$ is the depolarizing error probability. Therefore, for a Markovian noise model we have $\frac{p^{(s)}}{3}=\frac{1}{4}\big{[}1-\exp(-\tau/T_{2})\big{]}$ [18] where $\tau$ is the time required to generate a photon. We note that a depolarizing channel, where $T_{2}=T_{1}$, is a worst-case assumption for the noise allowing us to infer a lower bound on the threshold for $T_{2}$ from $p^{(s)}$.
@@ VI.1 Logical Fusion with repetition codes
For synchronous FFCC, we seek to perform logical $\left\{XX,ZZ\right\}$ fusions on qubits encoded in a $Z$-repetition code of size $m$. In order to recover the $\overline{XX}$ outcome, all $m$ physical $XX$ outcome must be obtained (Eq. (2)), but for $\overline{ZZ}$ a single physical $ZZ$ outcome is sufficient. As such, it is beneficial to perform rotated physical fusions on each pair of physical qubits, which measure $XX$ and $ZZ$ on success, and recover the $XX$ outcome upon fusion failure (Sec. II.1). In this manner, the $\overline{XX}$ outcome is protected by the choice of physical fusion, while $\overline{ZZ}$ is protected by redundancy of the repetition code.
The probabilities of errors and erasures on fusion outcomes using a repetition code with fixed code size (REP) are calculated analytically in Supplementary Note II. We employ a simple majority-voting decoding strategy for errors on the $k$ non-erased $ZZ$ outcomes, such that $\left\lfloor\frac{k-1}{2}\right\rfloor$ errors can be corrected. The $Z$-repetition code offers no tolerance to $Z$ error, so the error rate on the $\overline{XX}$ outcome, provided it is not erased, is the probability of having an odd number of erroneous $XX$ outcomes.
@@ VI.2 Repeat-until-success logical fusions
The probabilities for different logical fusion outcomes of a single RUS fusion under photon loss and errors can be analytically derived, see Supplementary Note II. Using these analytical expressions, logical fusion outcomes are sampled to obtain a syndrome graph for calculating fault-tolerant thresholds.
@@ VII Fault-tolerant Thresholds
The loss-tolerance thresholds are presented in Fig. 7a, for various code sizes $m$ for REP and maximal number of physical fusion attempts $N$ for RUS. The highest REP loss threshold obtained is $\ell=2.29\%$ for $m=5$. For RUS logical fusion, we observe that $\ell$ improves significantly as the number of maximally allowed fusion attempts $N$ increases, achieving $\ell=7.4\%$ with $N=10$. Using a strategy where additional $N$ re-attempts of RUS are allowed with both emitters re-initialized from the previous layer of logical fusions (Supplementary Note III), $\ell$ can reach up to $8\%$.
@@
Fig. 8a plots the simulated spin $Z$ error threshold with the number of fusion attempts per logical fusion, with $p^{(s)}_{X}=p^{(s)}_{Y}=0$. We observe that for spin errors occurred during the generation of encoded linear chains, the RUS threshold increases up to $p^{(s)}_{Z}=0.59\%$ at $N=10$ whereas for REP it peaks at $p^{(s)}_{Z}=0.19\%$ then steadily decreases to $0.095\%$. As discussed in Sec. V.2, a spin $Z$ error will localize as a $Z$ error in the next emitted photon. For REP the number of fusion attempts $m$ is fixed, such that when $m$ grows the probability of flipping $\overline{XX}$ also increases, and this leads to the decreasing threshold for high $m$. In the limit of low $m$, however, more fusion attempts increases the fusion success probability resulting in a more connected and hence more robust lattice, with an optimal threshold at $m=5$. For RUS, most logical fusions terminate in a few attempts assuming no photon loss (Supplementary Note II), so that only $Z$ errors occurring in the photons emitted during those attempts contribute to a flip in $\overline{XX}$. As a consequence, the spin $Z$ error threshold saturates as we increase $N$.
For spin depolarizing error (Fig. 8b), both thresholds again increase for low number of fusion attempts thanks to a better connectivity of the graph as explained above. The RUS (REP) threshold peaks at $p^{(s)}=0.2\%$ ($0.125\%$) then gradually decreases. This can be understood by first considering the propagation of spin $X$/$Y$ errors during resource state generation into photon errors, before mapping these photon errors to logical fusion errors. A spin $X$/$Y$ error occurred while generating the first $k\leq m$ photons in a logical qubit will lead to $X/Y$ errors in the subsequently generated $m-k$ photons  [^1] More precisely, a spin $Y$ error results in a $Y$ error on the first emitted photon and $X$ errors on the next generated photons in the current logical qubit $i$, as well as a $Z$ error on the first photon of the next qubit $i+1$., as well as a $Z$ photon error on the first photon of the next logical qubit (Sec. V.2). This implies that the probability of flipping the $ZZ$ outcome (hence $\overline{ZZ}$) increases with the number of fusion attempts, since a larger $m$ results in a greater probability of $X/Y$ spin errors within an $m$-photon logical qubit. Both error thresholds therefore reduce at higher values of $m$/$N$. Moreover, for REP, majority voting becomes ineffective at correcting erroneous $ZZ$ outcome due to a high weight of $X/Y$ errors from a single spin error.